OpenStudy (anonymous):

four men and four women line up at a checkout counter in a store. In how many different ways can they line up if the first person in line is a man, and the people in line alternate man, woman, man, woman, and so on?

OpenStudy (anonymous):

counting principle all the way for these problems

OpenStudy (anonymous):

how many choices for the first person in line?

OpenStudy (anonymous):

if it is not clear, let me know and we can walk through it slowly (not too slowly)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

whew ok good then how many for the second person?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

right now we go faster then 3 then 3, then 2 then 2, then 1 then 1 counting principle says to multiply all this mess together you can ignore the 1

OpenStudy (anonymous):

soooo 576?

OpenStudy (anonymous):

final answer is \[4\times 4\times 3\times 3\times 2\times 2\]or if you prefer \[4!\times 4!\]

OpenStudy (anonymous):

okay, thanks

OpenStudy (anonymous):

yeah, 576

OpenStudy (anonymous):

nice tumblr btw